OpenStudy (anonymous):

Approximate to the nearest tenth the real zeros of g(x)=2x^5+3x-2

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Newton Rhapson method?

OpenStudy (anonymous):

what's that?

OpenStudy (anonymous):

i never met that method? never heard of it

terenzreignz (terenzreignz):

It's the only method I know to approximate zeroes :(

OpenStudy (anonymous):

is there a formula?

terenzreignz (terenzreignz):

If you've never heard of it, chances are it's not the correct way to solve this...

OpenStudy (anonymous):

ahaha

OpenStudy (anonymous):

do you know other methods in getting the real zeros

terenzreignz (terenzreignz):

Nope :(

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

i'' ll pick another one

OpenStudy (anonymous):

solve for b in the equation\[\sqrt{3^{b}}=2^{b +1}\]

terenzreignz (terenzreignz):

First things first... you have to express them with the same base, right? Which one do you want, 2 or 3?

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

Okay, so... let's do this... \[\huge 3^{\frac12b}=2^{b+1}\]

terenzreignz (terenzreignz):

welp... actually, it might be easier to just get the natural logarithm of both sides... \[\huge \ln3^{\frac12b}=\ln2^{b+1}\]

terenzreignz (terenzreignz):

See where it goes from here?

OpenStudy (anonymous):

what does ln stands for?

OpenStudy (anonymous):

is that the same with log

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

It's the logarithm with base e, or approximately 2.718

OpenStudy (anonymous):

ahhhhhhhh

terenzreignz (terenzreignz):

So, due to the properties of logarithms, we get... \[\huge \frac{1}{2}b\ln \ 3 = (b+1)\ln \ 2\] and solve.

OpenStudy (anonymous):

i don't know ? ? ?

terenzreignz (terenzreignz):

ln 3 and ln 2 are just constants. You can double both sides, giving you \[\huge b\ln \ 3 = 2(b+1)\ln \ 2\]

OpenStudy (anonymous):

you multiply both sides with 2 ?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

I just did. To make it simpler, at least. Now let's rearrange a bit\[\huge b\ln \ 3 = (b+1)2\ln \ 2\]

terenzreignz (terenzreignz):

Properties of logarithms again, to make it more pleasing to the eye... \[\huge b\ln \ 3 = (b+1)\ln \ 2^2\]

terenzreignz (terenzreignz):

So finally b ln 3 = (b+1) ln 4 Now solve.

OpenStudy (anonymous):

will i divide both sides by\[\ln 3\]

terenzreignz (terenzreignz):

It's a start. \[\huge b =(b+1)\frac{\ln \ 4}{\ln \ 3}\]

terenzreignz (terenzreignz):

For simplicity, let's let \[\huge k = \frac{\ln \ 4}{\ln \ 3}\]

OpenStudy (anonymous):

k stand for CONSTANT, isn't it?

OpenStudy (anonymous):

but where does the b+1

terenzreignz (terenzreignz):

Just so that it's easier. Now solve for b b = (b+1)k

OpenStudy (anonymous):

\[\frac{ b }{b +1 }=k\]

terenzreignz (terenzreignz):

We're not solving for k, we're solving for b.

OpenStudy (anonymous):

\[\frac{ b }{ b +1}=\ln 4\div \ln 3\]

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

It's right, but it doesn't help you, it's b you want to solve for.

OpenStudy (anonymous):

what's the right thing to do? i'm already comfused?

terenzreignz (terenzreignz):

b = (b+1)k Distribute k b = bk + k then bring bk to the left b - bk = k factor out b b(1-k) = k divide both sides by k \[\huge b = \frac{k}{1-k}\]

OpenStudy (anonymous):

will i substitute now the value f k ?

OpenStudy (anonymous):

will i substitute now the value f k ?\[b =\frac{\frac{ \ln 4 }{\ln 3 }}{1+\frac{ \ln 4 }{\ln 3 }}\]

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

0.557885891

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

don't have a calculator. as long as you keyed in the correct values, you should be fine.

OpenStudy (anonymous):

TJ i had an answer but this answer doesn't fit with it

OpenStudy (anonymous):

our teacher had checked it. . . my answer was right but i don't have the solutions. . now, our teacher needed the solutions as our project

OpenStudy (anonymous):

the answer there is -4.8188

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Problem... I honestly don't know :/

OpenStudy (anonymous):

?????????

terenzreignz (terenzreignz):

I don't know how to get to your answer :(

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Wait, I know now... You put a plus sign instead of a minus sign.... Review the posts earlier :P

OpenStudy (anonymous):

hahahaa i see the difference

terenzreignz (terenzreignz):

Good.

OpenStudy (anonymous):

-4.818841679

OpenStudy (anonymous):

i reallly got the answer yepey . . thanks or your help . . . TJ

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

do you know somebody who's good in statistics

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

No... Just post in a new question, and start with the words "STATISTICS QUESTION" That ought to attract anybody good in Stat :)

OpenStudy (anonymous):

can you convince her or him to help me ??????please

OpenStudy (anonymous):

ok thanks 4 the advice see you tomorrow bye good night MY KNIGHT AND SHINING ARMOR

terenzreignz (terenzreignz):

:)